Circular No. 8053

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002kj Further to IAUC 8027, W. M. Wood-Vasey, G. Aldering, P. Nugent,

and K. Li report the discovery of an apparent supernova (mag 18.5) on unfiltered NEAT images (cf. IAUC 8029) taken with the Haleakala 1.2-m telescope on 2002 Dec. 1.35, 1.36, and 1.37 UT. Confirming images were obtained with the same telescope on Dec. 17, showing the new object at mag 18.9. SN 2002kj is located at α = 0^{h}35^{m}24^{s}.73, δ = +3^{o}59′11′′.4 (equinox 2000.0), which is 1′′.6 west of the center of the apparent host galaxy. No point source is visible at this location in previous NEAT images taken on 2001 Aug. 14, Sept. 17, Oct. 6, Nov. 9, 2002 Oct. 28, 29, and Nov. 2, to limiting mag 21.0 (S/N = 3).

COMETS C/2001 RX_{14} (LINEAR) AND C/2002 V1 (NEAT) M. Honda, University of Tokyo, and colleagues (T. Yamashita, H.

Kataza, T. Miyata, T. Fujiyoshi, S. Sako, Y. K. Okamoto, T. Onaka, T. Sekiguchi, D. Kinoshita, and J. Watanabe) report on mid-infrared observations of two comets with the 8.2-m Subaru Telescope (+ COMICS). N/Qband imaging of C/2002 V1 on Jan. 11.2 UT yield the following total fluxes within a 2′′.73-box aperture: 8.8 μm, 0.83 ± 0.01 Jy; 11.7 μm, 1.55 ± 0.02 Jy; 12.4 μm, 1.95 ± 0.04 Jy; 18.8 μm, 2.79 ± 0.06 Jy. Low-resolution (250) N-band spectroscopic observations (range 8–13 μm) of C/2002 V1 showed a broad amorphous silicate feature with a 11.2-μm local peak, indicating the presence of crystalline olivine. Total fluxes for C/2001 RX_{14} on Jan. 11.6 (measured as above): 8.8 μm, 0.066 ± 0.004 Jy; 12.4 μm, 0.279 ± 0.016 Jy; 18.8 μm, 0.356 ± 0.022 Jy. Low-resolution N-band spectroscopy of C/2001 RX_{14} showed the possible shallow silicate feature, but no 11.2-μm local peak was found within the error bars.

COMET P/2003 A1

Elliptical orbital elements from MPEC 2003-A86:

T = 2003 Feb. 1.990 TT  ω = 357.^{o}421
e = 0.48182  Ω = 55.224 }2000.0
q = 1.91415 AU  i = 46.213

a = 3.69402 AU n^{o} = 0.138821 P = 7.10 years

2003 January 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green